a stone is rotated in a vertical circle circle speed at bottom most point is √8gR,where R is the radius of a circle ratio of the Tension at the top and the bottom is?

Grade:11

1 Answers

Khimraj
6 years ago
At bottom most point tension will be Tb = mg +mv2/R
So Tb =  mg + m(8gR)/R = 9mg
Let velocity at top most point is v then it is given by work energy theorem
Change in K.E. = Work Done
m(8gR)/2 – mv2/2 = mg(2R)
So mv2 = 4mgR
Tension at top most point is given by Tt = mv2/R -mg
then Tt = 4mgR/R – mg = 3mg
then Tt/Tb= 3mg/9mg = 1/3.
If still you have doubt then please clearify it. If you like answer then please approve it.
Thank YOU.